# Pumping and Steady Streaming driven by Two-Frequency Oscillations of a Cylinder

### v1

@@ 4.4. Net Force
Integrating the surface stress (32) over the cylinder, the net horizontal force is
@@
This force calculation gives insight as to why pumping does not occur for the single-frequency case. From Section 4.2.1, the only steady term though third order in amplitude is proportional to $\sin(2\theta)$, and the only nonzero terms in the pressure expansion, equation (31), are proportional to $\cos(2\theta)$ and $\cos(4\theta)$. Therefore there is no net force for single-frequency oscillations through third order in amplitude, though we have not yet considered the possibility of a net force resulting at higher order for the single-frequency case. In the next section, we analyze the form of the solution at higher orders, and we extend the analysis to general frequency ratio to derive necessary conditions on the frequency ratio for pumping.
@@ 5. Analysis of General Frequency Ratios
In the previous section we showed that the steady stream function for frequency ratio $\alpha=2$ contained a term at third order proportional to $\sin(\theta)$ which indicates there is a net force on the fluid. In this section we examine the solution structure of $\psi_{n}$, the solution to the $n^{th}$ order perturbative equation (38), to determine whether there is a net force on the fluid, and, if so, at what order, for general frequency ratios.
For this analysis it is convenient to express the motion of the cylinder as
@@
where gcd(a,b)=1(a,b)=1. This expression is a slight modification from equation (1), and it amounts to different definition of the Reynolds number than the one used in the previous sections. We again change reference frames by fixing the cylinder, and the governing equations have the form,
@@
but the number of nonzero terms is finite at each order. The boundary condition determines the nonzero terms at leading order, and at higher orders the structure of the solution depends on the products of the lower order terms. In the following sections we first determine which terms proportional to $\sin\left(m\theta\right)$ are nonzero at each order, and then we derive conditions for the presence of a steady term ($k=0$).
@@ 5.1. $\theta$-dependent Terms
The form of the right-hand side of (38) dictates the solution structure at higher orders. Substituting $\psi_{1}$ into the quadratic nonlinearity, the right-hand side of (38) is proportional to $\sin(\theta)\cos(\theta)$, and thus the is solution is proportional to $\sin(2\theta)$. Therefore, $\psi_{2}$ is of the form
@@ 5.2.1. No Pumping for Single Frequency
We demonstrate that single-frequency oscillations do not pump by showing that $\psi_{n}$ cannot have a steady component when $n$ is odd. We examine the frequencies that arise at each order and deduce the orders at which the steady terms occur. For example, Figure 5 of Willis & Hohenegger (2024) illustrates how the frequencies at each order arise from the lower order frequencies for the single-frequency case.
@@ 5.3. Numerical Results for Other Frequency Ratios
According to Result 5.3, a third order scaling between flux and amplitude is the smallest possible order for pumping and occurs only for frequency ratio $2$. Furthermore, Result 5.1 implies that pumping is an odd-ordered effect, and thus the next smallest order at which pumping can occur is fifth order. From Result 5.3, a fifth order scaling can only occur when the frequency ratio is $3/2$ or $4$, respectively, corresponding to $a=2,b=3$ or $a=1,b=4$. Figure 5(a) shows the fluxes from numerical simulations for different frequency ratios and amplutides, and these results verify pumping is fifth order in amplitude for frequency ratios $3/2$ and $4$. Moreover, the fluxes for frequency ratios $3/2$ and $4$ are significantly smaller than the fluxes for frequency ratio $2$. For example, at amplitude $0.7$, the flux of frequency ratio $2$ is approximately $41$ times larger than the flux at frequency ratio $3/2$. In Figure 5(b) we show the passive tracer particles over $10$ periods, and the net flow is not obvious without zooming in closely or playing movie 3 from the supplementary materials. Though the fluxes for frequency ratios $3/2$ and $4$ are small, they show clear scaling with amplitude. By contrast, the small fluxes for frequency ratio $3$ are not correlated with the amplitude. According to Result 5.3, there is no pumping for frequency ratio $3$, and the small fluxes reported in Figure 5(a) arise from numerical error.
@@
The general theory from Section 5 stems from a regular perturbation analysis, where we assume low Reynolds number so that there is no boundary layer. Simulation results presented in Section 3 for frequency ratio $2$ were performed at Re = 10. However, the numerical simulations for other frequency ratios were performed at Re = 40, where there is a boundary layer (see Figure 5(b)), because, at lower Reynolds numbers, the fluxes are small, and it is difficult to measure pumping at higher order. However, as discussed in Chong et al. (2013), the perturbative solution of Holtsmark et al. (1954) contains the large Reynolds number results of Schlichting (1932) and Wang (1968). Thus, the theory is still valid under larger Reynolds numbers, where a single boundary layer is present.
@@ 6. Discussion
Using asymptotic analysis and numerical simulations, we examined the flow around a cylinder moving with two-frequency oscillations, and we derived necessary conditions for pumping. Strikingly, the necessary conditions on the frequency ratio for non-zero net motion are identical to those observed in the physical system of objects sliding via frictional forces on a laterally vibrating surface (Reznik & Canny, 2001; Hashemi et al., 2022). In Hashemi et al. (2022), they identified the time asymmetry of the driving oscillation necessary for net motion. Specifically, the sliding object only exhibited net motion when the vibrations were non-antiperiodic. A $2\pi$-periodic function $f(t)$ is antiperiodic if there exists $0<\phi<2\pi$ such that
@@
Conversely, $f(t)$ is non-antiperiodic if $f(t)$ does not satisfy Equation (46) for any $\phi$. Hashemi et al. (2022) showed that for frequency ratio $\alpha=b/a$, the two-frequency oscillation in equation (1) is antiperiodic if both $a$ and $b$ are odd. Otherwise, the motion is non-antiperiodic; i.e. one of $a$ or $b$ is odd and the other is even. For example, for $\alpha=1/1$ and $\alpha=3/1$, the motion is antiperiodic (Equation (46) is satisfied for $\phi=\pi$), and for $\alpha=2/1$ the motion is non-anitperiodic.
@@
In this paper, we not only derived the necessary conditions for pumping but also obtained the smallest order in amplitude at which pumping occurs. Frequency ratio two results in pumping at third order in amplitude, and all other frequency ratios produce pumping at higher order. This result indicates that pumping is strongest for frequency ratio two at small amplitude. In numerical simulations, we observed (Figure 5(a)) that the flux for $\alpha=2$ is at least $10$ times larger than the flux from other frequency ratios. These results are consistent with those of Hashemi et al. (2022) who reported both experimentally and numerically that $\alpha=2$ generated the largest net motion. Likewise, Reznik & Canny (2001) and Hui et al. (2024) showed in their models that the largest net frictional force was produced when the frequency ratio was $\alpha=2$.
In our analysis and computational experiments, the two oscillations were of equal amplitude and in phase. Analysis and experiments on sliding frictional systems have considered the more general motion of the form
@@
For example, it was predicted that for frequency ratio two, the maximum sliding velocity occurs for $A_{2}/A_{1}\approx 0.25$ (Zhang et al., 2024a; Reznik & Canny, 1998), which was validated experimentally (Hui et al., 2024). The analysis from this work focused on necessary conditions on the frequency ratio for pumping based on the solution structure, and it does not predict how the pumping rate depends on parameters. The analytic solution can be obtained through second order (Holtsmark et al., 1954), but solving the resulting equations third order and higher, where pumping occurs, is not feasible.
@@
Steady streaming has been exploited in many microfluidic applications such as drug delivery (Sumner et al., 2021), particle trapping (Agarwal et al., 2018; Mutlu et al., 2018; Patel et al., 2014; Thameem et al., 2016; Volk et al., 2020; Wang et al., 2012a; Zhang et al., 2024b), bubble-driven flow (Marmottant et al., 2006; Rallabandi et al., 2014; Wang et al., 2013), mixing (Ahmed et al., 2009; Huang et al., 2013; Kumar et al., 2011; Liu et al., 2003), and pumping (Huang et al., 2014; Marmottant & Hilgenfeldt, 2004; Tovar & Lee, 2009; Tovar et al., 2011; Zhang & Rallabandi, 2024). Microchannel pumps driven by single-frequency steady streaming rely on design asymmetry of the apparatus to produce a net-directed flow. For example, Tovar & Lee (2009) designed a micropump in which cavities of trapped air bubbles were placed at an angle along the channel walls. The angling of the cavities produces an asymmetric streaming flow that results in a net force that drive the flow through the channel. In our work, we have shown that a cylinder vibrating with two-frequency oscillations can pump fluid in either direction. The asymmetry in this case is temporal rather than spatial, and this idea could lead to new designs of microfluidic pumps or potentially enhance the performance of existing designs.
@@
The autonomous propulsion of microparticles is a means of transporting objects by steady streaming (Ahmed et al., 2016; Collis et al., 2017; Li et al., 2024; Lippera et al., 2019; Nadal & Lauga, 2014; Nadal & Michelin, 2020; Sabrina et al., 2018; Wang et al., 2012b). The particle or the background flow vibrates with single-frequency oscillations, and variation in particle shape or density leads to an asymmetric streaming flow that induces a net propulsive force on the particle. The asymptotic analysis of this phenomenon involves computing how the particle asymmetry leads to a net force on itself (Nadal & Lauga, 2014; Collis et al., 2017). Our analysis illustrates how a net force can arise from temporal asymmetry through two-frequency oscillations of a symmetric object. Though our analysis was for a cylinder, it could be extended to a sphere. Based on our results, we predict that symmetric particles could be transported by employing two-frequency oscillations.
@@ Appendix A Spatial Structure of the Solution
We show prove Result 5.1 in Section 5 by proving a more general case involving multifrequency oscillations. Suppose the cylinder is oscillating in the horizontal direction,
@@
Then, fixing the cylinder, the governing equations for multifrequency oscillation are
$\left(\Delta-\text{Re}\frac{\partial}{\partial t}\right)\Delta\psi=\epsilon\text{Re}\;\boldsymbol{u}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi\right),$
@@
$\Delta\left(\Delta-\text{Re}\;\frac{\partial}{\partial t}\right)\psi_{n}=\sum_{i+j=n}\text{Re}\;\boldsymbol{u}_{i}\boldsymbol{\cdot}\boldsymbol{\nabla}(\Delta\psi_{j}),$ (48)
@@
To prove Result 5.1 we prove two lemmas in the following subsections that show that terms proportional to $\sin(\theta)$ and $\cos(\theta)$ only appear at odd order in the expansions for the velocity and pressure, respectively. In Section 4.4 we showed that a net force requires such terms, which together with the below lemmas prove Result 5.1.
@@ A.1. Solution Structure of $\psi_{n}$
$\text{Re}\;\boldsymbol{u}_{1}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi_{1}\right)=\frac{\text{Re}}{2}\left(\frac{1}{r}F_{1}^{1}(r,t)\cdot\frac{\partial}{\partial r}D_{1}F_{1}^{1}(r,t)-\frac{1}{r}\frac{\partial}{\partial r}F_{1}^{1}(r,t)\cdot D_{1}F_{1}^{1}(r,t)\right)\sin(2\theta),$
@@
$\begin{array}[]{ll}\boldsymbol{u}_{i}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi_{j}\right)&=\displaystyle\sum_{s=1}^{i/2}\sum_{q=0}^{(j-1)/2}\frac{2s}{r}F_{2s}^{i}(r,t)\frac{\partial}{\partial r}D_{2q+1}F_{2q+1}^{j}(r,t)\sin((2q+1)\theta)\cos(2s\theta)\\[10.0pt] \displaystyle&-\displaystyle\sum_{s=1}^{i/2}\sum_{q=0}^{(j-1)/2}\frac{(2q+1)}{r}D_{2q+1}F_{2q+1}^{j}(r,t)\frac{\partial}{\partial r}F_{2s}^{i}(r,t)\sin(2s\theta)\cos((2q+1)\theta),\end{array}$ (55)

### v2

@@ 4.4. Net Force
Integrating the surface stress (34) over the cylinder, all the terms corresponding to $k\geq 2$ vanish due to symmetry, and the net horizontal force is
@@
This force calculation gives insight as to why pumping does not occur for the single-frequency case. From Section 4.2.1, the only steady term through third order in amplitude is proportional to $\sin(2\theta)$, and the only nonzero term in the pressure expansion, equation (33), is proportional to $\cos(2\theta)$. Therefore there is no net force for single-frequency oscillations through third order in amplitude, though we have not yet considered the possibility of a net force resulting at higher order for the single-frequency case. In the next section, we analyze the form of the solution at higher orders, and we extend the analysis to general frequency ratio to derive necessary conditions on the frequency ratio for pumping.
@@ 5. Analysis of General Frequency Ratios
In the previous section, we showed that the steady stream function for frequency ratio $\alpha=2$ contained a term at third order proportional to $\sin(\theta)$. In this section, we examine the solution structure of $\psi_{n}$, the solution to the $n^{th}$ order perturbative equation (40), to determine whether there is potentially a net force on the fluid due to the presence of a $\sin(\theta)$ term in the expansion and if so, what order.
For this analysis, it is convenient to express the motion of the cylinder as
@@
where gcd(a,b)=1(a,b)=1. This expression is a slight modification from equation (1), and it amounts to a different definition of the Reynolds number than the one used in the previous sections. We again change reference frames by fixing the cylinder, and the governing equations have the form,
@@
but the number of nonzero terms is finite at each order. The boundary condition determines the nonzero terms at leading order, and at higher orders the structure of the solution depends on the products of the lower order terms. In the following sections, we first determine which terms proportional to $\sin\left(m\theta\right)$ are nonzero at each order, and then we derive conditions for the presence of a steady term ($k=0$).
@@ 5.1. $\theta$-dependent Terms
The form of the right-hand side of (40) dictates the solution structure at higher orders. Substituting $\psi_{1}$ into the quadratic nonlinearity, the right-hand side of (40) is proportional to $\sin(\theta)\cos(\theta)$, and thus the solution is proportional to $\sin(2\theta)$. Therefore, $\psi_{2}$ is of the form
@@ 5.2.1. No Pumping for Single Frequency
We demonstrate that single-frequency oscillations do not pump by showing that $\psi_{n}$ cannot have a steady component when $n$ is odd. We examine the frequencies that arise at each order and deduce the orders at which the steady terms occur. For example, Figure 5 of Willis and Hohenegger (2024) illustrates how the frequencies at each order arise from the lower order frequencies for the single-frequency case.
@@ 5.3. Numerical Results for Other Frequency Ratios
According to Result 5.3, for small amplitude $\epsilon$, the strength of pumping scales like $\epsilon^{n}$, where $n=a+b$ is the sum of the two frequencies and must be odd. Thus, the strongest pumping results when $n=3$, corresponding to frequency ratio $\alpha=2$. The next smallest order at which pumping can occur is fifth order. From Result 5.3, a fifth order scaling can only occur when the frequency ratio is $3/2$ or $4$, respectively, corresponding to $a=2,b=3$ or $a=1,b=4$. Figure 7(a) shows the fluxes from numerical simulations for different frequency ratios and amplitudes, and these results verify pumping is fifth order in amplitude for frequency ratios $3/2$ and $4$. Moreover, the fluxes for frequency ratios $3/2$ and $4$ are significantly smaller than the fluxes for frequency ratio $2$. For example, at amplitude $0.7$, the flux of frequency ratio $2$ is approximately $41$ times larger than the flux at frequency ratio $3/2$. In Figure 7(b) we show the passive tracer particles over $10$ periods, and the net flow is not obvious without zooming in closely or playing movie 3 from the supplementary materials. Though the fluxes for frequency ratios $3/2$ and $4$ are small, they show clear scaling with amplitude. By contrast, the small fluxes for frequency ratio $3$ are not correlated with the amplitude. According to Result 5.3, there is no pumping for frequency ratio $3$, and the small fluxes reported in Figure 7(a) arise from numerical error. Note that in Section 3.2 we examined how the flux scaled with amplitude for $\alpha=2$ at Re = 10. Here, we use Re = 40 because at lower Reynolds numbers, the fluxes are small, and it is difficult to measure pumping at higher order.
@@
5.4. Near-Miss Frequencies
Figure 8. Time average of fluxes over one time unit for frequency ratios $\alpha$ = 2, $\alpha$ = 1.999, and $\alpha$ = 1.99 at amplitude $\epsilon=0.7$, and Re = 10. The inset shows the time-averaged fluxes to time $200$. On short time scales the fluxes are similar, but on long time scales, the fluxes for $\alpha=1.99$ and $\alpha=1.999$ oscillate with amplitude similar to the steady flux for $\alpha=2$.
An interesting question is what happens for frequency ratios that are close to, but not equal to, $\alpha=2$. Figure 8 shows the fluxes averaged over one time unit for frequency ratios $\alpha$ = 2, $\alpha$ = 1.999, and $\alpha$ = 1.99 at amplitude $\epsilon=0.7$, and Re = 10. The fluxes for $\alpha=1.99$ and $\alpha=1.999$ oscillate on long time scales, with periods 100 and 1000, respectively, with amplitude similar to the steady flux for $\alpha=2$. According to Result 5.3, for $\alpha=1.99$ and $\alpha=1.999$ the lowest order in amplitude pumping could occur is 299 and 2999, respectively. Over long periods we expect negligible net flux for these “near-miss” frequency ratios, but on a short time scale they exhibit fluxes similar to $\alpha=2$.
@@ 6. Discussion
Using asymptotic analysis and numerical simulations, we examined the flow around a cylinder moving with two-frequency collinear oscillations, and we derived necessary conditions for pumping. Strikingly, the necessary conditions on the frequency ratio for non-zero net motion are identical to those observed in the physical system of objects sliding via frictional forces on a laterally vibrating surface (Reznik and Canny, 2001; Hashemi et al., 2022). In Hashemi et al. (2022), they identified the time asymmetry of the driving oscillation necessary for net motion. Specifically, the sliding object only exhibited net motion when the vibrations were non-antiperiodic. A $2\pi$-periodic function $f(t)$ is antiperiodic if there exists $0<\phi<2\pi$ such that
@@
Conversely, $f(t)$ is non-antiperiodic if $f(t)$ does not satisfy Equation (48) for any $\phi$. Hashemi et al. (2022) showed that for frequency ratio $\alpha=b/a$, the two-frequency oscillation in equation (1) is antiperiodic if both $a$ and $b$ are odd. Otherwise, the motion is non-antiperiodic; i.e. one of $a$ or $b$ is odd and the other is even. For example, for $\alpha=1/1$ and $\alpha=3/1$, the motion is antiperiodic (Equation (48) is satisfied for $\phi=\pi$), and for $\alpha=2/1$ the motion is non-antiperiodic.
@@
In this paper, we not only derived the necessary conditions for pumping but also obtained the smallest order in amplitude at which pumping occurs. Frequency ratio two results in pumping at third order in amplitude, and all other frequency ratios produce pumping at higher order. This result indicates that pumping is strongest for frequency ratio two at small amplitude. In numerical simulations, we observed (Figure 7(a)) that the flux for $\alpha=2$ is at least $10$ times larger than the flux from other frequency ratios. These results are consistent with those of Hashemi et al. (2022) who reported both experimentally and numerically that $\alpha=2$ generated the largest net motion. Likewise, Reznik and Canny (2001) and Hui et al. (2024) showed in their models that the largest net frictional force was produced when the frequency ratio was $\alpha=2$.
Some insight into the connection between the flow around an oscillating cylinder and the frictional sliding problem is provided by Hashemi et al. (2024) in which they analyze the movement of a particle in a tank of fluid undergoing two-frequency oscillations. The model involves linear and nonlinear drag forces induced by the relative motion of the particle and the surrounding fluid. They perform analysis in the small amplitude limit and show that only frequency ratio $\alpha=2$ produces a nonzero time-average velocity at first-order in amplitude. Like the streaming problem we analyze, other frequency ratios produce net movement at higher order in amplitude. Hashemi et al. (2024) argues that both models of particle in an oscillating tank and of an object sliding on frictional surface are related to the same reduced model of a forced mechanical system with nonlinear dissipation.
In our analysis and computational experiments, the two oscillations were of equal amplitude, in phase, and collinear. Analysis and experiments on sliding frictional systems have considered the more general motion of the form
@@
For example, it was predicted that for frequency ratio two, the maximum sliding velocity occurs for $A_{2}/A_{1}\approx 0.25$ (Zhang et al., 2024b; Reznik and Canny, 1998), which was validated experimentally (Hui et al., 2024). The analysis from this work focused on necessary conditions on the frequency ratio for pumping based on the solution structure, and it does not predict how the pumping rate depends on parameters. The analytic solution can be obtained through second order (Holtsmark et al., 1954), but solving the resulting equations at third order and higher, where pumping occurs, is not feasible. The pumping rate for different amplitudes, phases, and Reynolds numbers could be measured experimentally or numerically. We note that to our knowledge there have been no experiments which examine the pumping rate generated by a cylinder oscillating with two frequencies. Another interesting generalization of the motion involves non-collinear, two-frequency vibrations (i.e., oscillations with different frequencies in two different directions). It may be possible to extend the analysis of this work to derive necessary conditions for a temporal asymmetry, and hence a net-directed flow, but more work on this problem is needed.
@@
Steady streaming has been exploited in many microfluidic applications such as drug delivery (Sumner et al., 2021), particle trapping (Agarwal et al., 2018; Mutlu et al., 2018; Patel et al., 2014; Thameem et al., 2016; Volk et al., 2020; Wang et al., 2012a; Zhang et al., 2024a), bubble-driven flow (Marmottant et al., 2006; Rallabandi et al., 2014; Wang et al., 2013), mixing (Ahmed et al., 2009; Huang et al., 2013; Kumar et al., 2011; Liu et al., 2003), and pumping (Huang et al., 2014; Marmottant and Hilgenfeldt, 2004; Tovar and Lee, 2009; Tovar et al., 2011; Zhang and Rallabandi, 2024). Microchannel pumps driven by single-frequency steady streaming rely on design asymmetry of the apparatus to produce a net-directed flow. For example, Tovar and Lee (2009) designed a micropump in which cavities of trapped air bubbles were placed at an angle along the channel walls. The angling of the cavities produces an asymmetric streaming flow that results in a net force that drive the flow through the channel. In our work, we have shown that a cylinder vibrating with two-frequency oscillations can pump fluid in either direction. The asymmetry in this case is temporal rather than spatial, and this idea could lead to new designs of microfluidic pumps or potentially enhance the performance of existing designs.
@@
The autonomous propulsion of microparticles is a means of transporting objects by steady streaming (Ahmed et al., 2016; Collis et al., 2017; Li et al., 2024; Lippera et al., 2019; Nadal and Lauga, 2014; Nadal and Michelin, 2020; Sabrina et al., 2018; Wang et al., 2012b). The particle or the background flow vibrates with single-frequency oscillations, and variation in particle shape or density leads to an asymmetric streaming flow that induces a net propulsive force on the particle. The asymptotic analysis of this phenomenon involves computing how the particle asymmetry leads to a net force on itself (Nadal and Lauga, 2014; Collis et al., 2017). Our analysis illustrates how a net force can arise from temporal asymmetry through two-frequency oscillations of a symmetric object. Though our analysis was for a cylinder, it could be extended to a sphere. Based on our results, we predict that symmetric particles could be transported by employing two-frequency oscillations.
@@ Appendix A Spatial Structure of the Solution
We prove Result 5.1 in Section 5 by proving a more general case involving multifrequency oscillations. Suppose the cylinder is oscillating in the horizontal direction,
@@
Then fixing the cylinder, the governing equations for multifrequency oscillation are
$\left(\Delta-Re\frac{\partial}{\partial t}\right)\Delta\psi=\epsilon Re\;\boldsymbol{u}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi\right),$
@@
$\Delta\left(\Delta-Re\;\frac{\partial}{\partial t}\right)\psi_{n}=\sum_{i+j=n}Re\;\boldsymbol{u_{i}}\boldsymbol{\cdot}\boldsymbol{\nabla}(\Delta\psi_{j}),$ (50)
@@
To prove Result 5.1, we prove two lemmas in the following subsections that show that terms proportional to $\sin(\theta)$ and $\cos(\theta)$ only appear at odd order in the expansions for the velocity and pressure, respectively. In Section 4.4 we showed that a net force requires such terms, which together with the below lemmas prove Result 5.1.
@@ A.1. Solution Structure of $\psi_{n}$
$Re\;\boldsymbol{u_{1}}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi_{1}\right)=\frac{Re}{2}\left(\frac{1}{r}F_{1}^{1}(r,t)\cdot\frac{\partial}{\partial r}D_{1}F_{1}^{1}(r,t)-\frac{1}{r}\frac{\partial}{\partial r}F_{1}^{1}(r,t)\cdot D_{1}F_{1}^{1}(r,t)\right)\sin(2\theta),$
@@
$\begin{array}[]{ll}\boldsymbol{u_{i}}\boldsymbol{\cdot}\boldsymbol{\nabla}\left(\Delta\psi_{j}\right)&=\displaystyle\sum_{s=1}^{i/2}\sum_{q=0}^{(j-1)/2}\frac{2s}{r}F_{2s}^{i}(r,t)\frac{\partial}{\partial r}D_{2q+1}F_{2q+1}^{j}(r,t)\sin((2q+1)\theta)\cos(2s\theta)\\[10.0pt] \displaystyle&-\sum_{s=1}^{i/2}\sum_{q=0}^{(j-1)/2}\frac{(2q+1)}{r}D_{2q+1}F_{2q+1}^{j}(r,t)\frac{\partial}{\partial r}F_{2s}^{i}(r,t)\sin(2s\theta)\cos((2q+1)\theta),\end{array}$ (57)